6 2/9 divided by 5 1/9 - 7/8 equals what?
Answer by checkley75(3666)

(6 2/9)/(5 1/9)-7/8
56/9*9/46-7/8
56/46-7/8
29/23-7/8=71/184 ANS.
IF YOU MEAN:
(6 2/9)/(5 1/9-7/8)
THEN:
(56/9)/(46/9-7/8)
(56/9)/(305/72)
56/9*72/305=56*8/305=448/305 ANS.